OpenStudy (lizzie.tttt):

What is the slope of a line that passes through the points, (3,1) and (-2,5) A; -4/5 B; 4/5 C; 5/4 D; -5/4

OpenStudy (aliqanber):

slope = (5-1) / (-2-3) = -4/5 or (1-5)/(3--2) = -4/5 it doesn't matter.